# A mathematical theory of topological invariants of quantum spin systems

### v1

@@ 3.1 Observables and derivations
We use the $\ell^{\infty}$ metric on ${\mathbb{R}}^{n}$, i. e. $d(x,y):=\max_{i=1,\ldots,n}|x_{i}-y_{i}|$. For any $U,V\subset{\mathbb{R}}^{n}$ we write $\operatorname{diam}(U):=\sup_{x,y\in U}d(x,y)$ and $d(U,V):=\inf_{x\in U,y\in V}d(x,y)$. They take values in extended non-negative reals $[0,\infty]$. Then $\operatorname{diam}(\varnothing)=0$ and $d(U,\varnothing)=\infty$ for any $U\subset{\mathbb{R}}^{n}$. For a nonempty set $U$ and $r\geq 0$ we define $U^{r}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq r\}$ while we set $\varnothing^{r}=\varnothing$.
A quantum lattice system consists of a subset $\Lambda\subset{\mathbb{R}}^{n}$ (”the lattice”) and a finite-dimensional complex Hilbert space $V_{j}$ for every $j\in\Lambda$. We make the following assumption on the lattice system [^6] In [11] $\Lambda$ was assumed to be a Delone set, i.e. it was also required to be uniformly filling. This assumption was imposed on physical grounds and is not needed for any of the results proved in [11]. In [8] $\Lambda$ was taken to be ${\mathbb{Z}}^{n}$ for simplicity [^7] Although we will need these assumptions in later sections, all of the results in Section 3.1 remain valid if we only assume that ${\mathscr{A}}(X)$ is finite-dimensional for any bounded $X\subset{\mathbb{R}}^{n}$.: there is a $C_{\Lambda}>0$ such that the number of points of $\Lambda$ in any hypercube of diameter $d$ is bounded by $C_{\Lambda}(d+1)^{n}$.
For any bounded nonempty $X\subset{\mathbb{R}}^{n}$ let ${\mathscr{A}}(X):=\bigotimes_{i\in X\cap\Lambda}\mathrm{Hom}_{{\mathbb{C}}}(V_{j},V_{j})$. For any $X\subset Y$ there is an inclusion ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ and the algebras ${\mathscr{A}}(X)$ form a direct system with respect to these inclusions. We extend this direct system to include the empty set by setting ${\mathscr{A}}(\varnothing)={\mathbb{C}}$ and letting the inclusion ${\mathscr{A}}(\varnothing)\hookrightarrow{\mathscr{A}}(X)$ take $\alpha\mapsto\alpha\boldsymbol{1}$. Each ${\mathscr{A}}(X)$ is a finite-dimensional C-algebra with the operator norm, and the inclusions ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ preserve this norm. The normed *-algebra of local observables is
@@
For any bounded $X\subset{\mathbb{R}}^{n}$, define the normalized trace ${\overline{\rm tr}}:{\mathscr{A}}(X)\to{\mathbb{C}}$ as ${\overline{\rm tr}}({\mathcal{A}})=\operatorname{tr}({\mathcal{A}})/{\sqrt{\dim({\mathscr{A}}(X))}}$. For any bounded $X\subset Y$ the partial trace ${\overline{\rm tr}}_{X^{c}}:{\mathscr{A}}(Y)\to{\mathscr{A}}(X)$ is uniquely specified by the condition ${\overline{\rm tr}}_{X^{c}}({\mathcal{A}}\otimes{\mathcal{B}})={\overline{\rm tr}}(A){\mathcal{B}}$ for any ${\mathcal{A}}\in{\mathscr{A}}(Y\backslash X)$ and ${\mathcal{B}}\in{\mathscr{A}}(X)$. Besides forming a direct system with respect to inclusions, the spaces ${\mathscr{A}}(X)$ are also an inverse system with respect to the partial trace. ${\overline{\rm tr}}$ extends to a normalized positive linear functional on ${\mathscr{A}}$, i.e. a state. We say that ${\mathcal{A}}\in{\mathscr{A}}$ is traceless if ${\overline{\rm tr}}({\mathcal{A}})=0$. The space of traceless anti-hermitian elements of ${\mathscr{A}}(X)$ will be denoted ${{\mathfrak{d}}_{l}}(X)$. ${{\mathfrak{d}}_{l}}(X)$ is a real Lie algebra with respect to the commutator. The Lie algebras ${{\mathfrak{d}}_{l}}(X)$ form a direct system over the directed set of bounded subsets of ${\mathbb{R}}^{n}$, and its limit will be denoted ${{\mathfrak{d}}_{l}}$. Equivalently, ${{\mathfrak{d}}_{l}}$ is the Lie algebra of traceless anti-hermitian elements of ${\mathscr{A}}_{\ell}$. Note that ${\mathscr{A}}_{\ell}={\mathbb{C}}{\mathbf{1}}\oplus({{\mathfrak{d}}_{l}}\otimes{\mathbb{C}})$.
@@
where $(k_{1},\ldots,k_{n})$, $(\ell_{1},\ldots,\ell_{n})$, and $(m_{1},\ldots,m_{j})$ are $n$-tuples of integers. We write ${\mathbb{B}}_{n}$ for the set of all bricks in ${\mathbb{R}}^{n}$.
The set of bricks exhaust the collection of bounded subsets of ${\mathbb{R}}^{n}$ in the sense that any bounded set is contained in a brick. In addition, the set of bricks satisfies the following regularity property:
@@
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,j))^{-2n-2}$ $\leq\frac{\pi^{4}(n+1)^{2}}{36}$ (5)
@@
Any pair of points ${\mathbf{x}},{\mathbf{y}}\in{\mathbb{Z}}^{n}$ specifies a brick with ${\mathbf{x}}$ and ${\mathbf{y}}$ on opposing corners, and any brick can be specified this way (not uniquely). With $X$ the brick corresponding to ${\mathbf{x}}$ and ${\mathbf{y}}$ it is easy to see that $\max(d({\mathbf{x}},j),d({\mathbf{y}},j))\leq\operatorname{diam}(X)+d(X,j)$, and so $(1+d({\mathbf{x}},j))(1+d({\mathbf{y}},j))\leq(1+\operatorname{diam}(X)+d(X,j))^{2}$. Thus we have
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,j))^{-2n-2}$ $\leq\sum_{\begin{subarray}{c}{\mathbf{x}},{\mathbf{y}}\in{\mathbb{Z}}^{n}\end{subarray}}(1+d({\mathbf{x}},j)^{-n-1}(1+d({\mathbf{y}},j)^{-n-1}$ (6)
$\leq\left(\sum_{{\mathbf{x}}\in{\mathbb{Z}}^{n}}(1+d({\mathbf{x}},j)^{-n-1}\right)^{2}$ (7)
and it remains only to bound the above sum. Let $f(k):=(1+k)^{-n-1}$ and $g(k):=\#({\mathbb{Z}}^{n}\cap B_{k}(j))\leq(1+k)^{n}$. Using summation by parts we have
$\sum_{j\in\Lambda}(1+d(x,j))^{-n-1}$ $\leq\sum_{k\geq 0}f(k)(g(k+1)-g(k))$ (8)
$=\lim_{k\to\infty}f(k)g(k)-\sum_{k\geq 0}g(k)(f(k+1)-f(k)).$ (9)
@@
We may use the partial trace to define the following space of operators which are intrinsically localized on a brick $Y$:
${{\mathfrak{d}}_{l}}^{Y}:=\{{\mathcal{A}}\in{{\mathfrak{d}}_{l}}(Y)\ |\ {\overline{\rm tr}}_{X^{c}}({\mathcal{A}})=0\text{ for any brick }X\subsetneq Y\}.$ (12)
This way, each ${{\mathfrak{d}}_{l}}(Y)$ decomposes as a direct sum ${{\mathfrak{d}}_{l}}(Y)=\bigoplus_{X\subseteq Y}{{\mathfrak{d}}_{l}}^{X}$ over bricks contained in $Y$, and for any brick $X\subseteq Y$ the partial trace ${\overline{\rm tr}}_{Y\backslash X^{c}}$ is the projection onto $\bigoplus_{Z\subseteq X}{{\mathfrak{d}}_{l}}^{Z}\subseteq{{\mathfrak{d}}_{l}}(Y)$.
@@
${\mathsf{F}}:{\mathcal{A}}\mapsto\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{Y},{\mathcal{A}}],$ (13)
@@
For any element ${\mathsf{F}}=\{{\mathsf{F}}^{Y}\}_{Y\in{\mathbb{B}}_{n}}$ of $\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ and every $U\subset{\mathbb{R}}^{n}$ write
@@
If $U$ is empty, then for $k>0$ $\|{\mathsf{F}}\|_{U,k}<\infty$ if and only if ${\mathsf{F}}^{Y}=0$ for all $Y\in{\mathbb{B}}_{n}$. Thus ${\mathfrak{D}}_{al}(\emptyset)=0$.
@@
$\|{\mathsf{F}}_{m}^{Y}-{\mathsf{F}}^{Y}\|$ $\leq\|{\mathsf{F}}_{m}^{Y}-{\mathsf{F}}_{N_{1}}^{Y}\|+\sum_{i=1}^{M-1}\|{\mathsf{F}}_{N_{i}}^{Y}-{\mathsf{F}}_{N_{i+1}}^{Y}\|+\|{\mathsf{F}}_{N_{M}}^{Y}-{\mathsf{F}}^{Y}\|$ (15)
$\leq\epsilon(1+\operatorname{diam}(Y)+d(U,Y))^{-k}+\|{\mathsf{F}}_{N_{M}}^{Y}-{\mathsf{F}}^{Y}\|.$ (16)
@@
In particular, ${\mathfrak{D}}_{al}(U)\subset{\mathfrak{D}}_{al}(V)$ and the inclusion is continuous.
@@
$1+\operatorname{diam}(Y)+d(Y,V)$ $\leq 1+\operatorname{diam}(Y)+d(Y,V^{r})+r$ (18)
$\leq 1+\operatorname{diam}(Y)+d(Y,U)+r$ (19)
$\leq(r+1)(1+\operatorname{diam}(Y)+d(Y,U)),$ (20)
@@
The above Lemma shows that the space ${\mathfrak{D}}_{al}(U)$ only depends on the asymptotic geometry of the region $U$ in the following sense: if $U\subseteq V^{r}$ and $V\subseteq U^{r}$ for some $r\geq 0$ then ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(V)$ (as subsets of $\prod_{Y}{{\mathfrak{d}}_{l}}^{Y})$ and are isomorphic as Fréchet spaces. In particular, for any non-empty bounded $U\subset{\mathbb{R}}^{n}$, the space ${\mathfrak{D}}_{al}(U)$ coincides with ${\mathfrak{D}}_{al}(\{0\})$ [^8] This is the space which in [7, 11] was denoted ${{\mathfrak{d}}_{al}}$.. To relate our approach to the traditional C-algebraic picture we prove the following:
@@
$\sum_{Y\in{\mathbb{B}}_{n}}\|{\mathsf{F}}^{Y}\|$ $\leq\|{\mathsf{F}}\|_{\{0\},2n+2}\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,0))^{-2n-2},$ (21)
and by Lemma 3.1 the above sum is finite. This shows that the map ${\mathfrak{D}}_{al}(\{0\})\to{\mathscr{A}}$ is well-defined and continuous. Its image is dense in the space of traceless anti-hermitian element of ${\mathscr{A}}$ because it contains the dense subspace ${{\mathfrak{d}}_{l}}$. Finally, to show that it is injective, writing ${\mathcal{A}}:=\sum_{Y\in{\mathbb{B}}_{n}}{\mathsf{F}}^{Y}$ we have
${\mathsf{F}}^{Y}={\overline{\rm tr}}_{Y^{c}}({\mathcal{A}})-\sum_{X\subsetneq Y}{\overline{\rm tr}}_{X^{c}}({\mathcal{A}})$ (22)
@@
We let ${{\mathfrak{d}}_{al}}\subset{\mathscr{A}}$ be the image of ${\mathfrak{D}}_{al}$ under the embedding of Prop. 3.2, with the Fréchet topology of ${\mathfrak{D}}_{al}$. We let ${\mathscr{A}}_{a\ell}={\mathbb{C}}{\mathbf{1}}\oplus({{\mathfrak{d}}_{al}}\otimes{\mathbb{C}})$.
@@
A derivation ${\mathsf{F}}\in{\mathfrak{D}}_{al}$ is inner iff it is contained in ${{\mathfrak{d}}_{al}}(U)$ for some bounded $U$.
Obviously, the Fréchet-Lie algebra of inner derivations can be identified with the Fréchet-Lie algebra ${{\mathfrak{d}}_{al}}$ of traceless anti-hermitian elements of ${\mathscr{A}}_{a\ell}$.
@@
The first inequality is always true – it follows from Lemma 3.2. For the second, let $Z\in{\mathbb{B}}_{n}$ and choose $x^{*},y^{*}\in Z$ so that $d(x^{*},U)=d(Z,U)$ and $d(y^{*},V)=d(Z,V)$. Then we have
$d(Z,U\cap V)$ $=\inf_{z\in Z}d(z,U\cap V)$ (25)
$\leq C\inf_{z\in Z}\max(d(z,U),d(z,V))$ (26)
$\leq C\inf_{z\in Z}\max(d(z,x^{*})+d(x^{*},U),d(z,y^{*})+d(y^{*},V))$ (27)
$\leq C(\operatorname{diam}(Z)+\max(d(Z,U),d(Z,V)),$ (28)
@@
$1+\operatorname{diam}(Z)+d(Z,U\cap V)\leq(C+1)(1+\operatorname{diam}(Z)+\max(d(Z,U),d(Z,V)),$ (29)
@@
    The sequence (30) admits a right splitting, i. e. a map $\gamma:{\mathfrak{D}}_{al}(U\cup V)\to{\mathfrak{D}}_{al}(U)\oplus{\mathfrak{D}}_{al}(V)$ with $\beta\circ\gamma=\operatorname{id}$. In particular, it is right-exact.
@@
    If $U$ and $V$ are transverse, then (30) is exact on the right.
@@
$\chi(Y):=\left\{\begin{array}[]{cc}1&\text{ if }d(Y,U)<d(Y,V)\\ 1/2&\text{ if }d(Y,U)=d(Y,V)\\ 0&\text{ if }d(Y,U)>d(Y,V)\end{array}\right..$ (34)
For any ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U\cup V)$ define $\gamma_{1}({\mathsf{F}}):=\sum_{Y\in{\mathbb{B}}_{n}}\chi(Y){\mathsf{F}}^{Y}$, and $\gamma_{2}({\mathsf{F}})={\mathsf{F}}-\gamma_{1}({\mathsf{F}})$. Then it is not hard to show that $\|\gamma_{1}({\mathsf{F}})\|_{U,k}\leq\|{\mathsf{F}}\|_{U\cup V,k}$ and $\|\gamma_{2}({\mathsf{F}})\|_{V,k}\leq\|{\mathsf{F}}\|_{U\cup V,k}$ and that $\gamma=(\gamma_{1},\gamma_{2})$ is a right-splitting of (30).
@@
$\pi_{i}(X\vee Y)$ $=\pi_{i}(X)\vee\pi_{i}(Y),$ (39)
$\operatorname{diam}(X)$ $=\max_{i=1,\ldots n}\operatorname{diam}(\pi_{i}(X)),$ (40)
$d(X,Y)$ $=\max_{i=1,\ldots n}d(\pi_{i}(X),\pi_{i}(Y)).$ (41)
@@
Let $X,Y,Z\in{\mathbb{B}}_{n}$ and let ${\mathsf{F}}\in{\mathscr{A}}(X)$ and ${\mathsf{G}}\in{\mathscr{A}}(Y)$. Then $[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}=0$ unless $X\cap Y\neq\varnothing$ and $Z\subseteq X\vee Y$.
@@
The requirement that $X\cap Y$ is clear, since ${\mathsf{F}}^{X}$ and ${\mathsf{G}}^{Y}$ would commute otherwise. Suppose $Z\nsubseteq X\vee Y$. Then $Z^{\prime}:=(X\vee Y)\cap Z$ is a brick that is strictly contained in $Z$, so $[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\in{\mathscr{A}}^{Z}\cap{\mathscr{A}}(Z^{\prime})=\{0\}.$
We make the following definintions for the next lemma. For any $U\subset{\mathbb{R}}^{n}$ and any brick $X$ write $\tilde{d}(X,U):=1+\operatorname{diam}(X)+d(X,U)$ and $\tilde{d}(X):=1+\operatorname{diam}(X)$.

### v2

@@ 3.1 Observables and derivations
We use the $\ell^{\infty}$ metric on ${\mathbb{R}}^{n}$, i. e. $d(x,y):=\max_{i=1,\ldots,n}|x_{i}-y_{i}|$. For any $U,V\subset{\mathbb{R}}^{n}$ we write $\operatorname{diam}(U):=\sup_{x,y\in U}d(x,y)$ and $d(U,V):=\inf_{x\in U,y\in V}d(x,y)$. They take values in extended non-negative reals $[0,\infty]$. Thus $\operatorname{diam}(\varnothing)=0$ and $d(U,\varnothing)=\infty$ for any $U\subset{\mathbb{R}}^{n}$. For a nonempty set $U$ and $r\geq 0$ we define $U^{r}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq r\}$ while we set $\varnothing^{r}=\varnothing$.
A quantum lattice system consists of a countable subset $\Lambda\subset{\mathbb{R}}^{n}$ (“the lattice”) and a finite-dimensional complex Hilbert space $V_{j}$ for every $j\in\Lambda$. We make the following assumption on the lattice system [^6] In [11] $\Lambda$ was assumed to be a Delone set, i.e. it was required to be uniformly filling and uniformly discrete. These assumptions were imposed on physical grounds. All the results proved in [11] hold under weaker assumptions adopted in this paper. In [8] $\Lambda$ was taken to be ${\mathbb{Z}}^{n}$ for simplicity. : there is a $C_{\Lambda}>0$ such that the number of points of $\Lambda$ in any hypercube of diameter $d$ is bounded by $C_{\Lambda}(d+1)^{n}$.
For any bounded nonempty $X\subset{\mathbb{R}}^{n}$ let ${\mathscr{A}}(X):=\bigotimes_{j\in X\cap\Lambda}\mathrm{Hom}_{{\mathbb{C}}}(V_{j},V_{j})$. For any $X\subset Y$ there is an inclusion ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ and the algebras ${\mathscr{A}}(X)$ form a direct system with respect to these inclusions. We extend this direct system to include the empty set by setting ${\mathscr{A}}(\varnothing)={\mathbb{C}}$ and letting the inclusion ${\mathscr{A}}(\varnothing)\hookrightarrow{\mathscr{A}}(X)$ take $\alpha\mapsto\alpha\boldsymbol{1}$. Each ${\mathscr{A}}(X)$ is a finite-dimensional $C^{*}$-algebra with the operator norm, and the inclusions ${\mathscr{A}}(X)\hookrightarrow{\mathscr{A}}(Y)$ preserve this norm. The normed *-algebra of local observables is
@@
For any bounded $X\subset{\mathbb{R}}^{n}$, define the normalized trace ${\overline{\rm tr}}:{\mathscr{A}}(X)\to{\mathbb{C}}$ as ${\overline{\rm tr}}({\mathcal{A}})=\operatorname{tr}({\mathcal{A}})/{\sqrt{\dim({\mathscr{A}}(X))}}$. For any bounded $X\subset Y$ the partial trace ${\overline{\rm tr}}_{X^{c}}:{\mathscr{A}}(Y)\to{\mathscr{A}}(X)$ is uniquely specified by the condition ${\overline{\rm tr}}_{X^{c}}({\mathcal{A}}\otimes{\mathcal{B}})={\overline{\rm tr}}({\mathcal{A}}){\mathcal{B}}$ for any ${\mathcal{A}}\in{\mathscr{A}}(Y\backslash X)$ and ${\mathcal{B}}\in{\mathscr{A}}(X)$. Besides forming a direct system with respect to inclusions, the spaces ${\mathscr{A}}(X)$ are also an inverse system with respect to the partial trace. ${\overline{\rm tr}}$ extends to a normalized positive linear functional on ${\mathscr{A}}$, i.e. a state. We say that ${\mathcal{A}}\in{\mathscr{A}}$ is traceless if ${\overline{\rm tr}}({\mathcal{A}})=0$. The space of traceless anti-hermitian elements of ${\mathscr{A}}(X)$ will be denoted ${{\mathfrak{d}}_{l}}(X)$. ${{\mathfrak{d}}_{l}}(X)$ is a real Lie algebra with respect to the commutator. The Lie algebras ${{\mathfrak{d}}_{l}}(X)$ form a direct system over the directed set of bounded subsets of ${\mathbb{R}}^{n}$, and its limit will be denoted ${{\mathfrak{d}}_{l}}$. Equivalently, ${{\mathfrak{d}}_{l}}$ is the Lie algebra of traceless anti-hermitian elements of ${\mathscr{A}}_{\ell}$. Note that ${\mathscr{A}}_{\ell}={\mathbb{C}}{\mathbf{1}}\oplus({{\mathfrak{d}}_{l}}\otimes{\mathbb{C}})$.
@@
where $(k_{1},\ldots,k_{n})$, $(\ell_{1},\ldots,\ell_{n})$, and $(m_{1},\ldots,m_{n})$ are $n$-tuples of integers. We write ${\mathbb{B}}_{n}$ for the set of all bricks in ${\mathbb{R}}^{n}$.
The set of bricks exhausts the collection of bounded subsets of ${\mathbb{R}}^{n}$ in the sense that any bounded subset is contained in a brick. In addition, the set of bricks satisfies the following regularity property:
@@
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,j))^{-2n-2}$ $\leq\frac{\pi^{4}4^{n}(n+1)^{2}}{36}$
@@
Any pair of points $x,y\in{\mathbb{Z}}^{n}$ specifies a brick with $x$ and $y$ on opposing corners, and any brick can be specified this way (not uniquely). With $X$ the brick corresponding to $x$ and $y$ it is easy to see that $\max(d(x,j),d(y,j))\leq\operatorname{diam}(X)+d(X,j)$, and so $(1+d(x,j))(1+d(y,j))\leq(1+\operatorname{diam}(X)+d(X,j))^{2}$. Thus we have
$\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,j))^{-2n-2}$ $\leq\sum_{\begin{subarray}{c}x,y\in{\mathbb{Z}}^{n}\end{subarray}}(1+d(x,j))^{-n-1}(1+d(y,j))^{-n-1}$
$=\left(\sum_{x\in{\mathbb{Z}}^{n}}(1+d(x,j))^{-n-1}\right)^{2},$
and it remains only to bound the above sum. Let $f(k):=(1+k)^{-n-1}$ and $g(k):=\#({\mathbb{Z}}^{n}\cap B_{k}(j))\leq(1+2k)^{n}$. Using summation by parts we have
$\sum_{j\in\Lambda}(1+d(x,j))^{-n-1}$ $\leq\sum_{k\geq 0}f(k)(g(k+1)-g(k))$
$=\lim_{k\to\infty}f(k)g(k)-\sum_{k\geq 0}g(k)(f(k+1)-f(k)).$
@@
For any brick $Y$ we define the following subspace of ${{\mathfrak{d}}_{l}}(Y)$:
${{\mathfrak{d}}_{l}}^{Y}:=\{{\mathcal{A}}\in{{\mathfrak{d}}_{l}}(Y)\ |\ {\overline{\rm tr}}_{X^{c}}({\mathcal{A}})=0\text{ for any brick }X\subsetneq Y\}.$
Each ${{\mathfrak{d}}_{l}}(Y)$ decomposes as a direct sum ${{\mathfrak{d}}_{l}}(Y)=\bigoplus_{X\subseteq Y}{{\mathfrak{d}}_{l}}^{X}$ over bricks contained in $Y$, and for any brick $X\subseteq Y$ the partial trace ${\overline{\rm tr}}_{Y\backslash X^{c}}$ is the projection onto $\bigoplus_{Z\subseteq X}{{\mathfrak{d}}_{l}}^{Z}\subseteq{{\mathfrak{d}}_{l}}(Y)$. Intuitively, ${{\mathfrak{d}}_{l}}^{Y}$ consists of elements of ${{\mathfrak{d}}_{l}}(Y)$ which are not localized on any brick properly contained in $Y$.
@@
${\mathsf{F}}:{\mathcal{A}}\mapsto\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{Y},{\mathcal{A}}],$
@@
For any element ${\mathsf{F}}=\{{\mathsf{F}}^{Y}\}_{Y\in{\mathbb{B}}_{n}}$ of $\prod_{Y\in{\mathbb{B}}_{n}}{{\mathfrak{d}}_{l}}^{Y}$ and every $U\subset{\mathbb{R}}^{n}$ we let
@@
If $U$ is empty, then for $k>0$ $\|{\mathsf{F}}\|_{U,k}<\infty$ if and only if ${\mathsf{F}}^{Y}=0$ for all $Y\in{\mathbb{B}}_{n}$. Thus ${\mathfrak{D}}_{al}(\emptyset)=0$. We also denote ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})={\mathfrak{D}}_{al}$.
@@
$\|{\mathsf{F}}_{m}^{Y}-{\mathsf{F}}^{Y}\|$ $\leq\|{\mathsf{F}}_{m}^{Y}-{\mathsf{F}}_{N_{1}}^{Y}\|+\sum_{i=1}^{M-1}\|{\mathsf{F}}_{N_{i}}^{Y}-{\mathsf{F}}_{N_{i+1}}^{Y}\|+\|{\mathsf{F}}_{N_{M}}^{Y}-{\mathsf{F}}^{Y}\|$
$\leq\epsilon(1+\operatorname{diam}(Y)+d(U,Y))^{-k}+\|{\mathsf{F}}_{N_{M}}^{Y}-{\mathsf{F}}^{Y}\|.$
@@
In particular, ${\mathfrak{D}}_{al}(U)\subseteq{\mathfrak{D}}_{al}(V)$ and the inclusion is continuous.
@@
$1+\operatorname{diam}(Y)+d(Y,V)$ $\leq 1+\operatorname{diam}(Y)+d(Y,V^{r})+r$
$\leq 1+\operatorname{diam}(Y)+d(Y,U)+r$
$\leq(r+1)(1+\operatorname{diam}(Y)+d(Y,U)),$
@@
The above Lemma shows that the space ${\mathfrak{D}}_{al}(U)$ only depends on the asymptotic geometry of the region $U$ in the following sense: if $U\subseteq V^{r}$ and $V\subseteq U^{r}$ for some $r\geq 0$ then ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(V)$ (as subsets of $\prod_{Y}{{\mathfrak{d}}_{l}}^{Y})$ and are isomorphic as Fréchet spaces. In particular, for any non-empty bounded $U\subset{\mathbb{R}}^{n}$, the space ${\mathfrak{D}}_{al}(U)$ coincides with ${\mathfrak{D}}_{al}(\{0\})$.
To relate the spaces ${\mathfrak{D}}_{al}(U)$ to the traditional $C^{*}$-algebraic picture we prove the following:
@@
$\sum_{Y\in{\mathbb{B}}_{n}}\|{\mathsf{F}}^{Y}\|$ $\leq\|{\mathsf{F}}\|_{\{0\},2n+2}\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,0))^{-2n-2},$
and by Lemma 3.1 the above sum is finite. This shows that the map ${\mathfrak{D}}_{al}(\{0\})\to{\mathscr{A}}$ is well-defined and continuous. Its image is dense in the space of traceless anti-hermitian elements of ${\mathscr{A}}$ because it contains the dense subspace ${{\mathfrak{d}}_{l}}$. Finally, to show that it is injective, writing ${\mathcal{A}}:=\sum_{Y\in{\mathbb{B}}_{n}}{\mathsf{F}}^{Y}$ we have
${\mathsf{F}}^{Y}={\overline{\rm tr}}_{Y^{c}}({\mathcal{A}})-\sum_{X\subsetneq Y}{\overline{\rm tr}}_{X^{c}}({\mathcal{A}})$
@@
Definition 3.3.
We let ${{\mathfrak{d}}_{al}}\subset{\mathscr{A}}$ be the image of ${\mathfrak{D}}_{al}(\{0\})$ under the embedding of Prop. 3.2, with the Fréchet topology of ${\mathfrak{D}}_{al}(\{0\})$.
In [11], the algebra ${\mathscr{A}}_{a\ell}$ of almost-local operators was defined as a subspace of ${\mathscr{A}}$ where a countable family of norms similar to (6) takes finite values. Here we equivalently define ${\mathscr{A}}_{a\ell}$ as the set of elements of ${\mathscr{A}}$ whose traceless hermitian and anti-hermitian parts live in ${{\mathfrak{d}}_{al}}$, topologized as ${\mathscr{A}}_{a\ell}={\mathbb{C}}{\mathbf{1}}\oplus({{\mathfrak{d}}_{al}}\otimes{\mathbb{C}})$.
@@
An element ${\mathsf{F}}\in{\mathfrak{D}}_{al}$ is inner iff it is contained in ${\mathfrak{D}}_{al}(U)$ for some bounded $U$.
@@
The first inequality is true even without assuming transversality – it follows from Lemma 3.2. For the second, let $Z\in{\mathbb{B}}_{n}$ and choose $x^{*},y^{*}\in Z$ so that $d(x^{*},U)=d(Z,U)$ and $d(y^{*},V)=d(Z,V)$. Then we have
$d(Z,U\cap V)$ $=\inf_{z\in Z}d(z,U\cap V)$
$\leq C\inf_{z\in Z}\max(d(z,U),d(z,V))$
$\leq C\inf_{z\in Z}\max(d(z,x^{*})+d(x^{*},U),d(z,y^{*})+d(y^{*},V))$
$\leq C(\operatorname{diam}(Z)+\max(d(Z,U),d(Z,V)),$
@@
$1+\operatorname{diam}(Z)+d(Z,U\cap V)\leq$
$\hskip 28.45274pt\leq(C+1)(1+\operatorname{diam}(Z)+\max(d(Z,U),d(Z,V))),$ (9)
@@
    The sequence (10) admits a right splitting, i.e. a map $\gamma:{\mathfrak{D}}_{al}(U\cup V)\to{\mathfrak{D}}_{al}(U)\oplus{\mathfrak{D}}_{al}(V)$ with $\beta\circ\gamma=\operatorname{id}$. In particular, it is exact on the right.
@@
    If $U$ and $V$ are transverse, then (10) is exact on the left.
@@
$\chi(Y):=\left\{\begin{array}[]{cc}1&\text{ if }d(Y,U)<d(Y,V)\\ 1/2&\text{ if }d(Y,U)=d(Y,V)\\ 0&\text{ if }d(Y,U)>d(Y,V)\end{array}\right..$
For any ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U\cup V)$ define $\gamma_{1}({\mathsf{F}}):=\sum_{Y\in{\mathbb{B}}_{n}}\chi(Y){\mathsf{F}}^{Y}$, and $\gamma_{2}({\mathsf{F}})={\mathsf{F}}-\gamma_{1}({\mathsf{F}})$. Then it is not hard to show that $\|\gamma_{1}({\mathsf{F}})\|_{U,k}\leq\|{\mathsf{F}}\|_{U\cup V,k}$ and $\|\gamma_{2}({\mathsf{F}})\|_{V,k}\leq\|{\mathsf{F}}\|_{U\cup V,k}$ and that $\gamma=(\gamma_{1},\gamma_{2})$ is a right splitting of (10).
@@
$\pi_{i}(X\vee Y)$ $=\pi_{i}(X)\vee\pi_{i}(Y),$
$\operatorname{diam}(X)$ $=\max_{i=1,\ldots n}\operatorname{diam}(\pi_{i}(X)),$
$d(X,Y)$ $=\max_{i=1,\ldots n}d(\pi_{i}(X),\pi_{i}(Y)).$
@@
Let $X,Y,Z\in{\mathbb{B}}_{n}$ and let ${\mathsf{F}}\in{{\mathfrak{d}}_{l}}(X)$ and ${\mathsf{G}}\in{{\mathfrak{d}}_{l}}(Y)$. Then $[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}=0$ unless $X\cap Y\neq\varnothing$ and $Z\subseteq X\vee Y$.
@@
The requirement that $X\cap Y\neq\varnothing$ is clear, since ${\mathsf{F}}^{X}$ and ${\mathsf{G}}^{Y}$ would commute otherwise. Suppose $Z\nsubseteq X\vee Y$. Then $Z^{\prime}:=(X\vee Y)\cap Z$ is a brick that is strictly contained in $Z$, so $[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\in{{\mathfrak{d}}_{l}}^{Z}\cap{{\mathfrak{d}}_{l}}(Z^{\prime})=\{0\}.$
We make the following definitions for the next lemma. For any $U\subset{\mathbb{R}}^{n}$ and any brick $X$ write $\tilde{d}(X,U):=1+\operatorname{diam}(X)+d(X,U)$ and $\tilde{d}(X):=1+\operatorname{diam}(X)$.